OpenStudy (anonymous):

write an equation of the line that is perpendicular to the line y=-x+6 and passes through (-2,-4)

OpenStudy (anonymous):

okay, first you have to find out the perpendicular slope of y=-x+6 -x is the slope, so the perpendicular slope is the negative reciprocal so the negative slope for -x is x so the slope is positive 1 (another example would be from 4 to -1/4) then you plug in the points given, and the perpendicular slope into point slope form (y1-y2=m(x1-x2) **remember m is slope) next you get y--4=1(x--2) you can then simplify to y+4=x+2 and then subtract 4 from both sides getting y=x-2. Voila math! **keep in mind, you can leave in in point-slope form [ y--4=1(x--2) ] ut since its asking for a line, i suggest keeping it as y=x-2.**

OpenStudy (anonymous):

Thank you so much! (:

OpenStudy (anonymous):

I don't get the 4to -1/4 part, please help!

OpenStudy (anonymous):

oh, that was just an example of taking a slope, and getting the perpendicular slope. if you had the slop of 4, to get the perpendicular slope from it, you would take 4 (basically 4/1) and flip the numerator and denominator. it then becomes 1/4. then to finish making it perpendicular, you make it negative of what it was----so if it was negative, you make it positive-- if it was positive, you make it negative.---- so it would then become -1/4. this was but an example, it was in no part of the problem so no worries!

OpenStudy (anonymous):

Oh okay! :D But to do the (m=y2-y1/x2-x1) don't I need to have 2 set of points not just one?

OpenStudy (anonymous):

yes, to find the slope, you must have two coordinates